# Towards Differential Handling of Various Blur Regions for Accurate Image Deblurring

### v1

@@ Abstract
Image deblurring aims to restore high-quality images by removing undesired degradation. Although existing methods have yielded promising results, they either overlook the varying degrees of degradation across different regions of the blurred image, or they approximate nonlinear function properties by stacking numerous nonlinear activation functions. In this paper, we propose a differential handling network (DHNet) to perform differential processing for different blur regions. Specifically, we design a Volterra block (VBlock) to integrate the nonlinear characteristics into the deblurring network, avoiding the previous operation of stacking the number of nonlinear activation functions to map complex input-output relationships. To enable the model to adaptively address varying degradation degrees in blurred regions, we devise the degradation degree recognition expert module (DDRE). This module initially incorporates prior knowledge from a well-trained model to estimate spatially variable blur information. Consequently, the router can map the learned degradation representation and allocate weights to experts according to both the degree of degradation and the size of the regions. Comprehensive experimental results show that DHNet effectively surpasses state-of-the-art (SOTA) methods on both synthetic and real-world datasets.
@@ 1 Introduction
pmage deblurring seeks to restore high-quality images from these blur versions. Due to the ill-posed nature, traditional method [20] attempts to tackle it by imposing various priors to limit the solution space. However, creating such priors is difficult and often lacks generalizability, making them impractical for real-world applications.
@@ 2.2 Volterra Series
The Volterra series is a model for nonlinear behavior that effectively captures ”memory” effects [37]. With its capability to learn nonlinear functions, Volterra Filters have been applied in deep learning [33, 1, 46, 32]. To enhance non-linearity and complement traditional activation functions, [46] introduces a single layer of Volterra kernel-based convolutions followed by standard CNN layers. VNN [32] proposes a cascaded approach of Volterra Filtering to significantly reduce the number of parameters. VolterraNet [1] presents a novel higher-order Volterra convolutional neural network designed for data represented as samples of functions on Riemannian homogeneous spaces. MR-VNet [33] employs Volterra layers to optimally introduce non-linearities in the image restoration process. In this paper, we design a VBlock that approximates nonlinearity without the computational burden of stacking multiple nonlinear functions.
@@ 3.1 Overall Pipeline
Our proposed DHNet, depicted in Figure 5, employs a hierarchical encoder-decoder architecture to facilitate effective hierarchical representation learning. Each encoder-decoder includes a DHBlock, which consists of $N_{i\in[1,2,3,4,5,6,7,8,9]}$ VBlocks and a degradation degree recognition expert module (DDRE). Given a degraded image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, DHNet first applies convolution to extract shallow features $\mathbf{F_{s}}\in\mathbb{R}^{H\times W\times C}$ (where $H$, $W$, and $C$ represent the height, width, and number of channels of the feature map, respectively). These shallow features pass through a four-scale encoder sub-network, where the resolution is progressively reduced while the number of channels increases. The resulting in-depth features then move to a middle block, and the deepest features are processed by a four-scale decoder, which gradually restores them to their original size. Finally, we apply convolution to the refined features to generate a residual image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$. This residual image is added to the degraded image to produce the restored image: $\mathbf{\hat{I}}=\mathbf{X}+\mathbf{I}$. We optimize the proposed network DHNet by minimizing the following loss function:
$L_{c}$ $=\sqrt{||\hat{I}-\overline{I}||^{2}+\epsilon^{2}}$
$L_{f}$ $=||\mathcal{F}(\hat{I})-\mathcal{F}(\overline{I})||_{1}$
@@ 3.3 Volterra Block
The strong performance and versatility of deep learning-based image deblurring models [10, 27] can be attributed to their nature as universal approximators. They achieve this by stacking numerous nonlinear activation functions, allowing them to fit any nonlinear function and effectively map the complex relationships between inputs and outputs. In order to reduce the system complexity caused by an excess of nonlinear activation functions, we design a Volterra block (VBlock) that utilizes the Volterra kernel to explore non-linearity within the network. Instead of relying on traditional nonlinear activation functions, VBlock employs higher-order convolutions to enhance linear convolution by enabling interactions between image pixels. As shown in Figure 5, the VBlock takes an input tensor $F$,first applies layer normalization (LN), and then encodes channel-wise context by utilizing 1×1 convolutions followed by 3×3 depth-wise convolutions to obtain the feature $F_{1}$ as follows:
@@
where $W_{p}^{(\cdot)}$ denotes the $1\times 1$ point-wise convolution, and $W_{d}^{(\cdot)}$ represents the $3\times 3$ depth-wise convolution. The resulting feature $F_{1}$ is then passed through the second-order Volterra kernel to map complex input-output relationships by facilitating interactions between image pixels. The process is as follows:
@@
Proof. Any continuous nonlinear function can be approximated by a polynomial. Using a Taylor expansion at $x_{0}$, the nonlinear function $\sigma(\cdot)$ can be represented as:
@@ 4.3.3 Design Choices for DDRE
Table 5: Results of alternatives to DDRE. Where $S$ and $T$ denote the number of experts and routers, respectively.
@@ 5 Conclusion
In this paper, we propose a differential handling network (DHNet) aimed at accurately deblurring images by processing different blur regions with fewer nonlinear activation functions. Specifically, we design a DHBlock consisting of VBlock and DDRE. The VBlock utilizes the Volterra kernel to explore non-linearity within the network, eliminating the need to stack numerous nonlinear functions to map complex input-output relationships. Meanwhile, the DDRE integrates prior knowledge from a well-trained model to estimate spatially variable blur information, allowing the router to map the learned degradation representation and assign weights to experts based on the degree of degradation and the size of the regions. Experimental results demonstrate that our DHNet outperforms state-of-the-art approaches.
@@ 7 More Proof for VBlock
A Volterra kernel with $L$ terms, it can be expressed as:
@@ 8 Additional Visual Results
In this section, we present additional visual results alongside state-of-the-art methods to highlight the effectiveness of our proposed approach, as shown in Figures 13,14,15. It is clear that our model produces more visually appealing outputs for both synthetic and real-world motion deblurring compared to other methods.
Figure 13: Comparison of image motion deblurring on the GoPro dataset [29] among MRLPFNet [11], AdaRevD [27], Restormer [43], FSNet [10], MambaIR [18], ALGNet [16], and our DHNet.
Figure 15: Comparison of image motion deblurring on the RealBlur dataset [31] among MRLPFNet [11], AdaRevD [27], Restormer [43], FFTformer [22], MambaIR [18], ALGNet [16], and our DHNet.

### v3

@@ Abstract
Image deblurring aims to restore high-quality images by removing undesired degradation. Although existing methods have yielded promising results, they either overlook the varying degrees of degradation across different regions of the blurred image. In this paper, we propose a differential handling network (DHNet) to perform differential processing for different blur regions. Specifically, we design a Volterra block (VBlock) to incorporate nonlinear characteristics into the deblurring network, enabling it to map complex input-output relationships without relying on nonlinear activation functions. To enable the model to adaptively address varying degradation degrees in blurred regions, we devise the degradation degree recognition expert module (DDRE). This module initially incorporates prior knowledge from a well-trained model to estimate spatially variable blur information. Consequently, the router can map the learned degradation representation and allocate weights to experts according to both the degree of degradation and the size of the regions. Comprehensive experimental results show that DHNet effectively surpasses state-of-the-art (SOTA) methods on both synthetic and real-world datasets.
@@ 2.2 Volterra Series
The Volterra series is a model for nonlinear behavior that effectively captures ”memory” effects [37]. With its capability to learn nonlinear functions, Volterra Filters have been applied in deep learning [33, 1, 46, 32]. To enhance non-linearity and complement traditional activation functions, [46] introduces a single layer of Volterra kernel-based convolutions followed by standard CNN layers. VNNs [32] proposes a cascaded approach of Volterra Filtering to significantly reduce the number of parameters. VolterraNet [1] presents a novel higher-order Volterra convolutional neural network designed for data represented as samples of functions on Riemannian homogeneous spaces. MR-VNet [33] uses Volterra layers to effectively introduce nonlinearities into the image restoration process. However, in MR-VNet, only the second-order nonlinearity is captured, while the first-order nonlinearity is overlooked. This results in reduced attention to local features and a higher risk of gradient explosion. In this paper, we design a VBlock that addresses these issues and approximates nonlinearity without the computational burden of stacking multiple nonlinear functions.
@@ 3.1 Overall Pipeline
Our proposed DHNet, depicted in Figure 3, employs a hierarchical encoder-decoder architecture to facilitate effective hierarchical representation learning. Each encoder-decoder includes a DHBlock, which consists of $N_{i\in[1,...,9]}$ VBlocks and a degradation degree recognition expert module (DDRE). Given a degraded image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, DHNet first applies convolution to extract shallow features $\mathbf{F_{s}}\in\mathbb{R}^{H\times W\times C}$ (where $H$, $W$, and $C$ represent the height, width, and number of channels of the feature map, respectively). These shallow features pass through a four-scale encoder sub-network, where the resolution is progressively reduced while the number of channels increases. The resulting in-depth features then move to a middle block, and the deepest features are processed by a four-scale decoder, which gradually restores them to their original size. Finally, we apply convolution to the refined features to generate a residual image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$. This residual image is added to the degraded image to produce the restored image: $\mathbf{\hat{I}}=\mathbf{X}+\mathbf{I}$.
@@ 3.3 Volterra Block
The strong performance and versatility of deep learning-based image deblurring models [10, 27] can be attributed to their nature as universal approximators. They achieve this by stacking numerous nonlinear activation functions, allowing them to fit any nonlinear function. In order to reduce the system complexity caused by an excess of nonlinear activation functions, we design a Volterra block (VBlock) that utilizes the Volterra kernel to explore non-linearity within the network. Instead of relying on traditional nonlinear activation functions, VBlock employs higher-order convolutions to enhance linear convolution by enabling interactions between image pixels. As shown in Figure 3, unlike MR-VNet [33], which treats the Volterra filter as a complete block, our VBlock incorporates it as part of a block. The VBlock first captures local features using LN, 1x1, and 3x3 convolutions before entering the Volterra filter, and then applies two 3x3 convolutions for each order branch. Specifically, our VBlock takes an input tensor $F$, first applies layer normalization (LN), and then encodes channel-wise context by utilizing 1×1 convolutions followed by 3×3 depth-wise convolutions to obtain the feature $F_{1}$ as follows:
@@
where $W_{p}^{(\cdot)}$ denotes the $1\times 1$ point-wise convolution, and $W_{d}^{(\cdot)}$ represents the $3\times 3$ depth-wise convolution.
The resulting feature $F_{1}$ is then processed by the second-order Volterra kernel to capture complex input-output relationships through interactions between image pixels. In this paper, we compute the second-order Volterra kernel as a product of traditional correlation operations. This representation can achieve an approximation error with arbitrary precision. Moreover, the separability assumption of the kernels enables efficient computation, which is especially advantageous in network settings. The process is as follows:
@@
Proof. Any continuous nonlinear function can be approximated by a polynomial. Similarly to VNNs [32], using a Taylor expansion at $x_{0}$, the nonlinear function $\sigma(\cdot)$ can be represented as:
@@ 4.3.3 Design Choices for DDRE
Table 5: Results of alternatives to DDRE, where $S$ and $T$ denote the number of experts and routers, respectively.
@@ 5 Conclusion
In this paper, we propose a differential handling network (DHNet) aimed at accurately deblurring images by processing different blur regions. Specifically, we design a DHBlock consisting of VBlock and DDRE. The VBlock utilizes the Volterra kernel to explore non-linearity within the network to map complex input-output relationships. Meanwhile, the DDRE integrates prior knowledge from a well-trained model to estimate spatially variable blur information, allowing the router to map the learned degradation representation and assign weights to experts based on the degree of degradation and the size of the regions. Experimental results demonstrate that our DHNet outperforms state-of-the-art approaches.
@@
6 Overview
The supplementary material is composed of:
Motivation Analysis: Sec.7
More Proof for VBlock: Sec.8
Loss Function: Sec.9
More Ablation Studies: Sec.10
Additional Visual Results: Sec.11
@@ 8 More Proof for VBlock
Our theoretical proof ideas are borrowed from VNNs [32]. A Volterra kernel with $L$ terms, it can be expressed as:
@@
Figure 11: Comparison of image motion deblurring on the GoPro dataset [29].
Figure 12: Comparison of image motion deblurring on the RealBlur dataset [31].
@@
9 Loss Function
To optimize the proposed network DHNet by minimizing the following loss function:
$L_{c}$ $=\sqrt{||\hat{I}-\overline{I}||^{2}+\epsilon^{2}}$
$L_{f}$ $=||\mathcal{F}(\hat{I})-\mathcal{F}(\overline{I})||_{1}$
10 More Ablation Studies
We provide more ablation studies on the GoPro dataset [29].
10.1 VBlock vs. MR-VNet [33]
Unlike MR-VNet [33], which treats the Volterra filter as a complete block, our VBlock incorporates it as part of a block. The VBlock first captures local features using LN, 1x1, and 3x3 convolutions before entering the Volterra filter, and then applies two 3x3 convolutions for each order branch. Additionally, by fusing the first- and second-order components and including a residual connection for the preorder, our VBlock effectively avoids the issues of local feature loss.
To validate the advantage of our VBlock over MR-VNet [33], we replace our VBlock with MR-VNet and present the results in Table 1. When replaced, the performance drops by 0.24 dB, demonstrating the superior performance of our VBlock. Additionally, MR-VNet [33] is prone to gradient collapse during practical training.
Table 7: Results of alternatives to VBlock.
Net MR-VNN [33] VBlock PSNR $\triangle$ PSNR
(a) 34.51 -
(b) 34.75 +0.24
10.2 Resource Efficient
We evaluate the model complexity of our proposed approach and other state-of-the-art methods in terms of running time and MACs. As shown in Table 8, our method achieves the lowest MACs value while delivering competitive performance in terms of running time.
Table 8: The evaluation of model computational complexity.
[…]
MPRNet [42] 1.148 777 32.66 0.959
Restormer [43] 1.218 140 32.92 0.961
FSNet [10] 0.362 111 33.29 0.963
MambaIR [18] 0.743 439 33.21 0.962
MR-VNet [33] 0.388 96 34.04 0.969
AdaRevD-L [27] 0.761 460 34.60 0.972
DHNet(Ours) 0.256 32 33.28 0.964
DHNet-B(Ours) 0.499 111 34.75 0.973
@@ 11 Additional Visual Results
In this section, we present additional visual results alongside state-of-the-art methods to highlight the effectiveness of our proposed approach, as shown in Figures 11,13,12. It is clear that our model produces more visually appealing outputs for both synthetic and real-world motion deblurring compared to other methods.